Solution A is 20% alcohol, and solution Bis 70% alcohol. How Much of each is needed to make50 liters of a solution that is 60% alcohol?
Answer by checkley77(12844)

.70x+.20(50-x)=.60*50
.70x+10-.20x=30
.50x=30-10
.50x=20
.50x=20
x=20/.50
x=40 liters of 70% alsohol is needed.
50-40=10 liters of 20% alcohol is needed.
Proof:
.70*40+.20*10=.60*50
28+2=30
30=30